OpenStudy (anonymous):

If x= 3 and r= 4, then what is y?

OpenStudy (lgbasallote):

\[y = \sqrt{r^2 - x^2}\] \[y = \sqrt{4^2 = 3^2}\]

OpenStudy (anonymous):

some other variable?

OpenStudy (anonymous):

man i must have missed that memo!

OpenStudy (lgbasallote):

nahh it's a divine seeing invisible thing

hero (hero):

\[4^2 \ne 3^2 \]

OpenStudy (lgbasallote):

uhh sorry that was a -

OpenStudy (lgbasallote):

\[y = \sqrt{4^2 - 3^2}\]